kindness has a son named gabriel. The sum of their ages is 84. After three years, kindness will be twice
as old as gabriel. What are their ages now?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

kindness has a son named gabriel. The sum of their ages is 84. After three years, kindness will be twice as old as gabriel. What are their ages now?
------------------
k + g = 84
k+3 = 2*(g+3)
----------------
Can you do the rest?

Answer by ikleyn(52803)

.
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problem for the day of April, 1
    - OVERVIEW of lessons on age problems
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".